# Deformation and breakup of the liquid ligament with various disturbances on the interface in shear flow

### v1

@@ 4.2 Analysis of the deformation of liquid ligament interface in shear flow
To investigate the effect of the flow field near the interface on the liquid ligament splitting, Fig. 25 shows the interface, velocity vectors, and streamline at $T=2.0$ for $\eta_{r}=0.9$ and $k_{r}=4.0$. The velocity vector and streamline are distributions in the $x$$z$ cross-section at $y/d=0$. As shown in Fig. 21(b), the nonlinear effect causes the liquid ligament to swell at $z/\lambda=0.15$, 0.3, 0.7, and 0.85. A large bulge also occurs at $z/\lambda=0.5$. A vortex due to the airflow is formed in the wake of each bulge. As these vortices induce a flow from the constricted part to the bulging part of the liquid ligament, the bulging part develops further, and the constricted part narrows further. As a result, the deformation of the ligament is accelerated by the vortex, and the ligament splits faster than the result of $\eta_{r}=0$ and $k_{r}=0$.
@@
Figure 26 shows the diameter of the breakup droplet. Table 1 shows the droplet diameter, which is the arithmetic average of all droplet diameters under each condition, and the standard deviation of the droplet diameter. In the case of $\eta_{r}=0$ and $k_{r}=0$, as the diameters of the droplets existing near $z/\lambda=0.4$ and 0.6 are much smaller than those of the droplets formed at $z/\lambda=0$ and 0.5, the average droplet diameter becomes small. On the other hand, the standard deviation is high due to the large variation in the diameters of the breakup droplets. The authors (Yanaoka and Nakayama, 2022) reported the droplet diameters of the main and satellite droplets, and the average diameter of the two droplets was $1.314d$. In this study, as the shear velocity is faster than that in the earlier study, the diameter of the droplet is smaller than that of the previous report. For $\eta_{r}=0.9$ and $k_{r}=2.0$, the trend is the same as for $\eta_{r}=0$ and $k_{r}=0$, and there is no significant difference in the average droplet size and its standard deviation. For $\eta_{r}=0.9$ and $k_{r}=4.0$, more droplets form than other conditions. Therefore, the total surface area of the liquid is the largest. Compared to the results of $\eta_{r}=0$, $k_{r}=0$ and $\eta_{r}=0.9$, $k_{r}=2.0$, the average droplet size is large, but the standard deviation is low. It can be seen from this that the droplet diameters were made uniform by adding the disturbance with the high wavenumber component. Therefore, for $\eta_{r}=0.9$ and $k_{r}=4.0$, it can be said that the atomization quality of the liquid ligament was improved.
We compare droplet diameters obtained in this study with existing studies (Rutland and Jameson, 1970; Lafrance, 1975; Lakdawala et al., 2015). A liquid jet generates the main droplet and satellite droplet. The diameter $d_{s}$ of the satellite droplet is smaller than the diameter $d_{m}$ of the main droplet. In this study, large and small droplets are also generated, similar to existing results. In Rutland and Jameson’s experiment (Rutland and Jameson, 1970), the diameters are $d_{m}/d=2.129$ and $d_{s}/d=1.127$ for $ka=0.34$. Lafrance (1975) reported $d_{m}/d=2.402$ and $d_{s}/d=1.450$ for $ka=0.32$. In Lakdawala et al.’s calculation (Lakdawala et al., 2015), the diameters are $d_{m}/d=2.350$ and $d_{s}/d=1.393$ for $ka=0.30$. In this study, we obtained smaller droplet diameters than existing results. In existing research (Rutland and Jameson, 1970; Lafrance, 1975; Lakdawala et al., 2015), the direction of shear velocity between gas and liquid is the same as the direction of the central axis direction of the liquid column. On the other hand, the two directions are orthogonal in this study. The physical properties of the liquid used also differ. It is considered that such a difference causes the difference in droplet diameters.
@@ 5 CONCLUSION
This study performed a numerical analysis of the deformation and breakup of a liquid ligament with various disturbances on the interface in shear flow. We investigated the effects of the airflow velocity difference, the wavenumber of the initial disturbance on the interface of the liquid ligament, and the high wavenumber disturbance on the deformation and splitting of the liquid ligament. As a result, the following findings were obtained.
@@
When an initial disturbance with a low wavenumber is applied to the interface, a new turbulence with a higher wavenumber than the initial disturbance is generated at the interface, and an intensive nonlinear effect appears. As the wavenumber of the disturbance applied to the interface increases, the liquid moving velocity along the central axis of the liquid ligament increases. In addition, on the high wavenumber side, the bulge of the liquid ligament develops earlier due to the vortex near the interface. As the wavenumber increases, the cross-sectional area of the constricted part of the liquid ligament decreases more quickly, and the breakup time of the liquid ligament becomes short.
In addition to the initial reference disturbance with a low wavenumber, when turbulence with twice the wavenumber of the reference disturbance is applied to the interface, the interface deformation and the splitting of the liquid ligament are similar to those with the single reference disturbance. Therefore, there is no significant difference between the average droplet size and its standard deviation under the two conditions. When turbulence with four times the wavenumber is added to the reference disturbance with the low wavenumber at the interface, the deformation of the liquid ligament is accelerated, and the liquid ligament splits faster. In addition, the total surface area of the liquid increases because the number of breakup droplets increases. The droplet diameters become uniform, and therefore, the atomization quality of the liquid ligament is improved.

### v2

@@ 4.2 Analysis of the deformation of liquid ligament interface in shear flow
To investigate the effect of the flow field near the interface on the liquid ligament splitting, Fig. 25 shows the interface, velocity vectors, and streamline at $T=2.0$ for $\eta_{r}=0.9$ and $k_{r}=4.0$. The velocity vector and streamline are distributions in the $x$$z$ cross-section at $y/d=0$. As shown in Fig. 21(b), the nonlinear effect causes the liquid ligament to swell at $z/\lambda=0.15$, 0.3, 0.7, and 0.85. A large bulge also occurs at $z/\lambda=0.5$. A vortex owing to the airflow is formed in the wake of each bulge. These vortices induce a flow from the constricted part of the liquid ligament to the bulging part; thus, the bulging part develops further, and the constricted part narrows further. Consequently, the deformation of the ligament is accelerated by the vortex, and the ligament splits faster than the result of $\eta_{r}=0$ and $k_{r}=0$.
@@
Figure 26 shows the diameter of the breakup droplet. Table 1 shows the droplet diameter, which is the arithmetic average of all droplet diameters under each condition, and the standard deviation of the droplet diameter. In the case of $\eta_{r}=0$ and $k_{r}=0$, the diameters of the droplets existing near $z/\lambda=0.4$ and 0.6 are much smaller than those of the droplets formed at $z/\lambda=0$ and 0.5; hence, the average droplet diameter becomes small. The standard deviation is high owing to the large variation in the diameters of the breakup droplets. The authors (Yanaoka and Nakayama, 2022) reported the droplet diameters of the main and satellite droplets, and the average diameter of the two droplets was $1.314d$. In this study, as the shear velocity is faster than that in the earlier study, the diameter of the droplet is smaller than that of the previous report. For $\eta_{r}=0.9$ and $k_{r}=2.0$, the trend is the same as for $\eta_{r}=0$ and $k_{r}=0$, and there is no significant difference in the average droplet size and its standard deviation. For $\eta_{r}=0.9$ and $k_{r}=4.0$, more droplets form than other conditions. Hence, the total surface area of the liquid is the largest. Compared to the results of $\eta_{r}=0$, $k_{r}=0$ and $\eta_{r}=0.9$, $k_{r}=2.0$, the average droplet size is large, but the standard deviation is low. It can be seen from this that the droplet diameters were made uniform by adding the disturbance with the high wavenumber component. Therefore, for $\eta_{r}=0.9$ and $k_{r}=4.0$, it can be said that the atomization quality of the liquid ligament was improved.
We compare droplet diameters obtained in this study with existing studies (Rutland and Jameson, 1970; Lafrance, 1975; Lakdawala et al., 2015). A liquid jet generates the main droplet and satellite droplet. The diameter $d_{s}$ of the satellite droplet is smaller than the diameter $d_{m}$ of the main droplet. In this study, large and small droplets are also generated, similarly to existing results. In Rutland and Jameson’s experiment (Rutland and Jameson, 1970), the diameters are $d_{m}/d=2.129$ and $d_{s}/d=1.127$ for $ka=0.34$. Lafrance (1975) reported $d_{m}/d=2.402$ and $d_{s}/d=1.450$ for $ka=0.32$. In Lakdawala et al.’s calculation (Lakdawala et al., 2015), the diameters are $d_{m}/d=2.350$ and $d_{s}/d=1.393$ for $ka=0.30$. In this study, we obtained smaller droplet diameters than existing results. In existing research (Rutland and Jameson, 1970; Lafrance, 1975; Lakdawala et al., 2015), the direction of shear velocity between gas and liquid is the same as the direction of the central axis direction of the liquid column. Conversely, the two directions are orthogonal in this study. The physical properties of the liquid used also differ. It is considered that such a difference causes the difference in droplet diameters.
@@ 5 CONCLUSION
This study performed a numerical analysis of the deformation and breakup of a liquid ligament with various disturbances on the interface in shear flow. We investigated the effects of the airflow velocity difference, the wavenumber of the initial disturbance on the interface of the liquid ligament, and the high wavenumber disturbance on the deformation and splitting of the liquid ligament. Consequently, the following findings were obtained.
@@
When an initial disturbance with a low wavenumber is applied to the interface, a new turbulence with a higher wavenumber than the initial disturbance is generated at the interface, and an intensive nonlinear effect appears. As the wavenumber of the disturbance applied to the interface increases, the liquid moving velocity along the central axis of the liquid ligament increases. Additionally, on the high wavenumber side, the bulge of the liquid ligament develops earlier owing to the vortex near the interface. As the wavenumber increases, the cross-sectional area of the constricted part of the liquid ligament decreases more quickly, and the breakup time of the liquid ligament becomes short.
In addition to the initial reference disturbance with a low wavenumber, when turbulence with twice the wavenumber of the reference disturbance is applied to the interface, the interface deformation and the splitting of the liquid ligament are similar to those with the single reference disturbance. Therefore, there is no significant difference between the average droplet size and its standard deviation under the two conditions. When turbulence with four times the wavenumber is added to the reference disturbance with the low wavenumber at the interface, the deformation of the liquid ligament is accelerated, and the liquid ligament splits faster. Additionally, the number of breakup droplets increases; hence, the total surface area of the liquid increases. The droplet diameters become uniform; therefore, the atomization quality of the liquid ligament is improved.